the problem is 14-(2q+5)= -2q+9
Answer by stanbon(48564)

14-(2q+5)= -2q+9
14-2q-5 = -2q+9
-2q+9=-2q+9
This equation will be true for
any value of q. Although it doesn't
look like it, the original problem
is an identity---true for all values
of the variable.
Cheers,
stan H.